# Automatically Detecting Online Deceptive Patterns

### v3

@@ 6.2 Findings
We also evaluate Shi et al.’s [64] DPGuard framework. To evaluate their framework, we first create a mapping between their taxonomy and the filtered taxonomy (see Section 11.7). For our evaluation, we only consider the categories mappable to the filtered taxonomy. Additionally, as the DPGuard framework does not provide localization capabilities, our evaluation is solely focused on DPGuard’s ability to identify deceptive patterns present within the webpage in our D3 Dataset. Our evaluation shows that DPGuard is unable to identify deceptive patterns with high accuracy, especially struggling to classify ‘hidden-subscription’ and ‘trick-wording’. These findings are consistent with the evaluation performed by Shi et al. [64].
@@ 7.1 Browser Extension for Web Users
The browser extension utilizes a hybrid architecture consisting of lightweight browser components with a locally running Flask server to run the AutoBot framework, similar to the Zotero extension [22]. On the browser side, we use a popup script to allow users to activate the extension. Once active, the extension takes a screenshot and sends it to the local Flask server for analysis. The Flask server hosts the AutoBot framework, using Flan-T5 as the language model. After analysis, the server returns the classifications to the extension, which are then rendered by the content script of the extension on the user’s screen. This rendering is shown in Figure 6.
@@
For the browser extension to be practical, it must perform its analysis in real time. We, therefore, conducted latency tests on its core Flan-T5 (Language) and YOLOv10 Ensemble (Vision) models using three different machine configurations representing various hardware capabilities (modern high-end with GPU, older mid-range, ARM-based) and comparing CPU versus GPU performance. Our results (detailed in Section 11.5) show that while performance on older hardware using only the CPU takes roughly 1.5 to 3 seconds per module (e.g., 1.95 seconds for YOLOv10 Ensemble on a CPU-only i5 laptop), newer devices with GPUs achieve near real-time performance. Specifically, on the 2023 laptop with a GPU, T5 inference (with 800 tokens as input) took less than 0.5 seconds, and YOLOv10 Ensemble performed its classifications in less than 0.3 seconds. Our experiments show the feasibility of implementing AutoBot as a browser extension, as modern hardware with GPU support enables a responsive experience without major performance delays [34, 4].
@@ 7.2 Lighthouse Reports for Web Developers
Research shows that while developers are often unaware of deceptive patterns on their websites and their impacts on users, they are open to addressing these patterns if properly informed [70]. To that end, we integrate AutoBot into Lighthouse [1, 29], a popular tool developers use to receive automated audits on website quality. Google bundles this tool with ChromeDev Tools, and major platforms such as Shopify, Wix, and Squarespace [^10] https://www.shopify.com/, https://www.wix.com/, https://www.squarespace.com/ integrate it into their workflows. We created a custom Lighthouse audit using AutoBot to fit directly into a web developer’s workflow. This audit uses an existing Lighthouse Gatherer [27] to capture screenshots of the website, processes it using AutoBot, and finally incorporates the findings into a Lighthouse report, as shown in Figure 7.
@@ 7.3 Enabling Web-Scale Analysis
We use this tool to analyze 11,118 websites, consisting of 6,626 diverse, popular domains selected from the Tranco list [54], and 4,492 e-commerce sites from the Shopify Partners Directory [^11] https://www.shopify.com/partners/directory. We detail our website selection process earlier in  Section 5.2.2.
@@ 8.1.1 Study Design
We conducted a within-subject study to assess the usability impact of highlighting deceptive patterns with AutoBot. We informed the participants that the objective of the survey was to test the usability of a web interface. Next, participants were asked to visit two custom-made websites, one with and one without highlighting, and fill out a System Usability Scale questionnaire [14] after each. In both these websites, the participants were asked to perform one of the four tasks: sign up, download, do shopping, or read a news article. Once the participants visit a website and hover over the highlighted text, a banner cautioning them about the deceptive pattern is shown. Note that the participants interacted with highlighted websites directly without installing a browser extension.
These custom-made websites were created by the authors based on various real websites, available at UXP2 Dark Patterns [^12] https://darkpatterns.uxp2.com/, caught using deceptive patterns. An example of such a website is shown in Section 11.8. After visiting each website, the participant is asked to complete a System Usability Scale questionnaire [14]. Additionally, we asked the participants to fill out a post-study questionnaire consisting of three questions:
@@ 8.2.4 Future Studies
The limited response rate suggests that a different approach is needed to interact with developers. While prior work had success with email-based notifications for web developers [70], those emails included language about vulnerabilities and legal liabilities. We opted not to use such language with the developers, which contributed to a reduced response rate. We suggest that future studies actively recruit developers and interview them about the Lighthouse reports. Another follow-up study can potentially help regulators generate variants of Lighthouse reports that check compliance with relevant laws, such as the French Data Protection Act.
@@ 9 Limitations
A limitation for AutoBot comes from the static UI analysis. Our approach inherently restricts the scope of detectable deceptive patterns to those visually present on a single page at a given time. Deceptive patterns such as nagging cannot be detected through this approach, and thus, we filter Gary et. al.’s taxonomy [31] to focus on static deceptive patterns.
@@ 10 Future Work
AutoBot is a framework that takes a screenshot and returns localized deceptive patterns. We provide three sample applications that build on top of it. We envision that applications like lighthouse-ci [28] can easily extend our work to integrate AutoBot into developer CI/CD workflows. Regulators can also use AutoBot with models aligned to their regulations to enforce policies automatically.
@@ 11.1 Full Taxonomy
Interference Confirmshaming Guilt-tripping users into making a specific choice “No, I prefer to pay more”
@@ 11.2 Website with Multiple Deceptive Pattern
Table 8: Mapping taxonomy of Deceptive Patterns to Gray et al. [31]

### v4

@@ 6.2 Findings
We also evaluate Shi et al.’s [50Shades] DPGuard framework. To evaluate their framework, we first create a mapping between their taxonomy and the filtered taxonomy (see Section 11.7). For our evaluation, we only consider the categories mappable to the filtered taxonomy. Additionally, as the DPGuard framework does not provide localization capabilities, our evaluation is solely focused on DPGuard’s ability to identify deceptive patterns present within the webpage in our D3 Dataset. Our evaluation shows that DPGuard is unable to identify deceptive patterns with high accuracy, especially struggling to classify ‘hidden-subscription’ and ‘trick-wording’. These findings are consistent with the evaluation performed by Shi et al. [50Shades].
@@ 7.1 Browser Extension for Web Users
The browser extension utilizes a hybrid architecture consisting of lightweight browser components with a locally running Flask server to run the AutoBot framework, similar to the Zotero extension [zotero]. On the browser side, we use a popup script to allow users to activate the extension. Once active, the extension takes a screenshot and sends it to the local Flask server for analysis. The Flask server hosts the AutoBot framework, using Flan-T5 as the language model. After analysis, the server returns the classifications to the extension, which are then rendered by the content script of the extension on the user’s screen. This rendering is shown in Figure 6.
@@
For the browser extension to be practical, it must perform its analysis in real time. We, therefore, conducted latency tests on its core Flan-T5 (Language) and YOLOv10 Ensemble (Vision) models using three different machine configurations representing various hardware capabilities (modern high-end with GPU, older mid-range, ARM-based) and comparing CPU versus GPU performance. Our results (detailed in Section 11.5) show that while performance on older hardware using only the CPU takes roughly 1.5 to 3 seconds per module (e.g., 1.95 seconds for YOLOv10 Ensemble on a CPU-only i5 laptop), newer devices with GPUs achieve near real-time performance. Specifically, on the 2023 laptop with a GPU, T5 inference (with 800 tokens as input) took less than 0.5 seconds, and YOLOv10 Ensemble performed its classifications in less than 0.3 seconds. Our experiments show the feasibility of implementing AutoBot as a browser extension, as modern hardware with GPU support enables a responsive experience without major performance delays [huggingface_transformersjs_webgpu, w3c_webgpu_2024].
@@ 7.2 Lighthouse Reports for Web Developers
Research shows that while developers are often unaware of deceptive patterns on their websites and their impacts on users, they are open to addressing these patterns if properly informed [stover_how_2023]. To that end, we integrate AutoBot into Lighthouse [lighthouse_chrome, lighthouse_overview], a popular tool developers use to receive automated audits on website quality. Google bundles this tool with ChromeDev Tools, and major platforms such as Shopify, Wix, and Squarespace [^10] https://www.shopify.com/, https://www.wix.com/, https://www.squarespace.com/ integrate it into their workflows. We created a custom Lighthouse audit using AutoBot to fit directly into a web developer’s workflow. This audit uses an existing Lighthouse Gatherer [lighthouse_gatherer] to capture screenshots of the website, processes it using AutoBot, and finally incorporates the findings into a Lighthouse report, as shown in Figure 7.
@@ 7.3 Enabling Web-Scale Analysis
We use this tool to analyze 11,118 websites, consisting of 6,626 diverse, popular domains selected from the Tranco list [pochat2018tranco], and 4,492 e-commerce sites from the Shopify Partners Directory [^11] https://www.shopify.com/partners/directory. We detail our website selection process earlier in Section 5.2.2.
@@ 8.1.1 Study Design
We conducted a within-subject study to assess the usability impact of highlighting deceptive patterns with AutoBot. We informed the participants that the objective of the survey was to test the usability of a web interface. Next, participants were asked to visit two custom-made websites, one with and one without highlighting, and fill out a System Usability Scale questionnaire [brooke1996sus] after each. In both these websites, the participants were asked to perform one of the four tasks: sign up, download, do shopping, or read a news article. Once the participants visit a website and hover over the highlighted text, a banner cautioning them about the deceptive pattern is shown. Note that the participants interacted with highlighted websites directly without installing a browser extension.
These custom-made websites were created by the authors based on various real websites, available at UXP2 Dark Patterns [^12] https://darkpatterns.uxp2.com/, caught using deceptive patterns. An example of such a website is shown in Section 11.8. After visiting each website, the participant is asked to complete a System Usability Scale questionnaire [brooke1996sus]. Additionally, we asked the participants to fill out a post-study questionnaire consisting of three questions:
@@ 8.2.4 Future Studies
The limited response rate suggests that a different approach is needed to interact with developers. While prior work had success with email-based notifications for web developers [stover_how_2023], those emails included language about vulnerabilities and legal liabilities. We opted not to use such language with the developers, which contributed to a reduced response rate. We suggest that future studies actively recruit developers and interview them about the Lighthouse reports. Another follow-up study can potentially help regulators generate variants of Lighthouse reports that check compliance with relevant laws, such as the French Data Protection Act.
@@ 9 Limitations
A limitation for AutoBot comes from the static UI analysis. Our approach inherently restricts the scope of detectable deceptive patterns to those visually present on a single page at a given time. Deceptive patterns such as nagging cannot be detected through this approach, and thus, we filter Gary et. al.’s taxonomy [gray2023towards] to focus on static deceptive patterns.
@@ 10 Future Work
AutoBotis a framework that takes a screenshot and returns localized deceptive patterns. We provide three sample applications that build on top of it. We envision that applications like lighthouse-ci [lighthouse-ci] can easily extend our work to integrate AutoBot into developer CI/CD workflows. Regulators can also use AutoBot with models aligned to their regulations to enforce policies automatically.
@@ 11 Conclusion
Acknowledgments
This work was supported by the NSF through awards CNS-1942014 and CNS-2247381, and by a research grant from the Google PSS Privacy Faculty Award program. Finally, we thank the reviewers for their thoughtful recommendations.
@@ 11.1 Full Taxonomy
Interface Interference Confirmshaming Guilt-tripping users into making a specific choice “No, I prefer to pay more”
@@ 11.2 Website with Multiple Deceptive Pattern
Table 8: Mapping taxonomy of Deceptive Patterns to Gray et al. [gray2023towards]